# Bosonization in $R$-paraparticle Luttinger models

### v2

@@ 5 Summary and Discussion
We have shown that bosonization applies to certain one-dimensional $R$-paraparticles systems described by the Luttinger model, to which flavor-charge separation occurs as a result. These $R$-paraparticles are classified as $R$-parafermions, which obey general exclusion principles and thus can only occupy a limited number of particle states. $R$-parafermionic density waves are shown to have bosonic signatures, but flavor waves generally do not exhibit this behavior. A comparison of the partition function in the Luttinger model shows that only $R$-parafermions that satisfy Pauli’s exclusion principle preserve their degeneracy when changing from the $R$-parafermion basis to the boson basis, further limiting the applicability of bosonization to $R$-paraparticles of order $p=1$. Nonetheless, these results show promises in potentially observing $R$-paraparticle signatures in one-dimensional systems.
We have motivated this study as a means of finding a general means of solving an $R$-paraparticle system in the presence of interaction. A dedicated section for such method was omitted in this paper since, by showing that a collection of $R$-parafermion operators can be recast as a boson operator, bosonization automatically implies that solutions applicable to fermionic Luttinger models should also apply to $R$-parafermionic systems (see Sec. 4.4C of [29]). Nonetheless, other generalizations such as the field-theoretic formalism of the Luttinger model should not be treated as trivial given the extensive use of generalized commutation relations in their definitions.
As mentioned in the previous section, generalization to arbitrary $m$ cases would benefit the development of possible experimental setups that may result to observations of $R$-paraparticle systems. This goes hand-in-hand with an understanding of the internal symmetries at play within the $R$-paraparticle. At the same time, a physical interpretation of a non-bosonic flavor wave oscillation is lacking, which may or may not affect potential observations of $R$-paraparticle singatures. Speculatively, this could also mean that $\hat{\sigma}_{r}(q)$ generally possess parabose statistics, but this remains an assumption until an example with such properties can be presented.
Acknowledgements
We thank M.A. Sulañgi for the helpful discussions.
Funding information
@@ Acknowledgements
D.F.S. acknowledges support from the DOST-SEI under the Accelerated Science and Technology Human Resource Development Program.
@@ Appendix A R-parabosons
In Sec. 2.2, we discussed the behavior of $R$-paraparticles where $d_{n>n^{\prime}}=0$ for some finite $n^{\prime}$. We now discuss what happens in the limit $n^{\prime}\to\infty$. We note that the single-mode partition functions of $R$-paraparticles satisfy [23]
$z_{R}(-x)z_{-R}(x)=1$ (A.1)
under the premise that $z_{R}(x)$ is interpreted as the Hilbert series [63] of $R$. Suppose $R$ describes an $R$-parafermion, then by Eq. (A.1),
$z_{-R}(x)=\frac{1}{\sum_{n=0}^{n^{\prime}}d_{n}(-x)^{n}}=\sum_{j=0}^{\infty}D_{j}x^{j},$ (A.2)
@@
$D_{0}=1,\quad D_{j}=\sum_{k=1}^{\min(j,n^{\prime})}d_{k}(-)^{k+1}D_{j-k}\in\mathbb{Z}_{+}.$ (A.3)
Eq. (A.2) applies only in the reduced domain $x\in(0,e^{-\beta\epsilon_{0}})$ under the following considerations:
  • By definition, $z_{R}(-x)=1+\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}$ ($d_{0}=0$ for all $R$-paraparticles, i.e., there is always only one vacuum state), so the second equality in Eq. (A.2) requires $1>\allowbreak|\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}|$. $1>-\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}$ is automatically satisfied by the fact that $z_{-R}(x)>\allowbreak 0$ as required from a partition function.
  • The minimum particle energy must be finite, i.e., $\epsilon_{0}\not\to$-$\infty$, as in this limit, $x\to\infty$, and
    $\lim_{x\to\infty}\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}=\begin{cases}+\infty,\quad\text{for even }n^{\prime}\\ -\infty,\quad\text{for odd }n^{\prime}\end{cases}.$ (A.4)
    The geometric expansion requires $\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}<1$, which is violated by the above equation, thus, $\epsilon_{0}$ must be finite.
  • Because $z_{-R}(e^{-\beta\epsilon_{0}})\not>0$, then $z_{-R}(x)$ must diverge at $x=e^{-\beta\epsilon_{0}}$. Note that $z_{-R}(e^{-\beta\epsilon_{0}})\neq 0$ because $z_{R}(-x)$ is a smooth function and therefore does not diverge for some finite $x$. This implies that $e^{-\beta\epsilon_{0}}$ is one of the roots of $z_{R}(-x)$.
From these, the average occupation number of R-R is therefore
$\langle\tilde{n}\rangle_{-R}=\frac{\sum_{n=1}^{n^{\prime}}nd_{n}(-)^{n-1}x^{n}}{\sum_{n=0}^{n^{\prime}}d_{n}(-x)^{n}}=\frac{x\partial_{-x}z_{R}(-x)}{z_{R}(-x)}.$ (A.5)
By Rolle’s theorem, any function will always have more real roots than their derivative, therefore, $z_{R}(-x)$ has at least one more zero than $\partial_{-x}z_{R}(-x)$, and thus $\langle\tilde{n}\rangle_{-R}$ will always be divergent in at least one $\epsilon$ value. There are no jump discontinuities in $\langle\tilde{n}\rangle_{-R}$, and removable discontinuities can always be factored out. Moreover,
$\lim_{\beta\epsilon\to\infty}\langle\tilde{n}\rangle_{-R}$ $=0,$ (A.6a)
@@
$\partial_{\beta\epsilon}\langle\tilde{n}\rangle_{-R}\Bigl{|}_{\beta\epsilon\to\infty}$ $\propto-d_{1},$ (A.6b)
@@
hence at the domain $\epsilon\in(\epsilon_{0},\infty)$, the average occupation number is exponentially decreasing from infinity, reminiscent of bose statistics. This motivates us to classify $R$-paraparticles that have single-mode partition functions in the form of Eq. (A.2) with $D_{n}\neq 0$ for all $n$ as $R$-parabosons.
@@
While all $R$-parafermions with a four-tensor $R$ have a corresponding $R$-parabosons described by R-R, the converse is not always true. Consider the examples
$M_{5}$ $=\begin{pmatrix}\alpha&0&0&0\\ 0&0&\gamma&0\\ 0&\gamma&0&0\\ \beta&0&0&-\alpha\end{pmatrix}$, (A.7a)
$M_{6}$ $=\begin{pmatrix}\alpha&0&0&\beta\\ 0&0&\gamma&0\\ 0&\gamma&0&0\\ 0&0&0&-\alpha\end{pmatrix}$, (A.7b)
@@
$M_{7}$ $=\begin{pmatrix}\alpha&0&0&0\\ 0&0&\delta&0\\ 0&1/\delta&0&0\\ 0&0&0&-\alpha\end{pmatrix}$, (A.7c)
where $\alpha,\gamma=\pm 1,\,\beta\in\mathbb{C},\,\delta\in\mathbb{C}_{\neq 0}$. In all cases, no exclusion principle can be extracted from their GCRs, and therefore all possible combinations of $\alpha,\beta,\gamma,$ and $\delta$ would result to $d_{n}\neq 0$ for all $n$. Moreover, their average occupation number
$\langle\tilde{n}\rangle=\frac{xz^{\prime}(x)}{z(x)}=\frac{\sum_{n=1}^{\infty}nd_{n}x^{n}}{\sum_{n=0}^{\infty}d_{n}x^{n}}$ (A.8)
will always diverge at the radius of convergence $x_{0}$ of $z(x)$. This is because $xz^{\prime}(x)$ grows faster than $z(x)$ as $x\to x_{0}^{-}$, and thus the resulting behavior of $\langle\tilde{n}\rangle$ is $R$-parabosonic. An implication of this result is that in the $R$-paraparticle formulation, there are more $R$-parabosons than $R$-parafermions.

### v3

@@ V Summary and Discussion
In this paper, we showed that $R$-paraparticles can be classified as $R$-parafermions or $R$-parabosons based on their occupation statistics. We worked mainly with $R$-parafermions, which have fermi-surface-like structures and possess generalized exclusion principles. We showed that bosonization applies to certain 1D $R$-parafermion systems described by the LM, in which flavor-charge separation occurs as a result. $R$-parafermionic density waves are shown to have bosonic signatures, however, flavor waves are bosonic in nature only for a subclass of $R$-parafermions. A comparison of the partition function in the LM shows that $R$-parafermions satisfy bosonization only at low temperatures. Nonetheless, these results show promise for potentially observing $R$-paraparticle signatures in 1D systems.
One of the main motivations of this study was to investigate a system of $R$-paraparticles in the presence of interaction. We have shown that for cases where bosonization is applicable, the treatment becomes closely similar to fermionic systems, speifically, their dispersion relations become equivalent. Nonetheless, other generalizations such as the field-theoretic formalism of the LM should not be treated as trivial given the extensive use of generalized commutation relations in their definitions.
Complication to our formalism arises when we consider $R$-paraparticle systems with more than two flavor degrees of freedom, as the current construction of the flavor wave operator makes sense only for $m=2$, in analogy to the known spin-$\frac{1}{2}$ LM. For a generalized theory, additional operators must be introduced which (i) preserves the number of degrees of freedom in the system under the basis transformation $\{\hat{\psi}_{r,k,a}\}_{a=1}^{m}\to\{\hat{\rho}_{r},\hat{\sigma}_{r,1},\ldots,\hat{\sigma}_{r,m-1}\}$, while at the same time (ii) allows for the description of systems even without prior knowledge of the nature of these degrees of freedom.
Beyond this issue, other possible extensions of this work would be to improve our understanding of the internal basis preference of $R$-paraparticles. Another open problem would be the interpretation of the requirement in Eq. (23), which, due to its role in bosonization for the LM, may be connected to the preferred internal basis of physically realizable $R$-parafermions. If not, then it begs the question of how non-bosonic flavor waves in certain $R$-parafermion systems should be interpreted. Furthermore, it is also interesting to explore the conditions under which emergent $R$-paraparticles appear in systems that are shown to host them.
Acknowledgements.
@@
D.F.S. acknowledges support from the DOST-SEI under the Accelerated Science and Technology Human Resource Development Program. We also acknowledge M.A. Sulañgi, Z. Wang, and F.N.C. Paraan for their helpful discussions and comments.
@@ Appendix A R-parabosons
In Sec. II.2, we discussed the behavior of paraparticles where $d_{n>n^{\prime}}=0$ for some finite $n^{\prime}$. We now discuss what happens in the limit $n^{\prime}\to\infty$.
To begin, we first note that the single-mode partition functions of paraparticles satisfy [32]
$z_{R}(-x)z_{-R}(x)=1$ (44)
under the premise that $z_{R}(x)$ is interpreted as the Hilbert series [62] of $R$. Suppose $R$ describes an $R$-parafermion, then by Eq. (44),
$z_{-R}(x)=\frac{1}{\sum_{n=0}^{n^{\prime}}d_{n}(-x)^{n}}=\sum_{j=0}^{\infty}D_{j}x^{j},$ (45)
@@
$D_{0}=1,\quad D_{j}=\sum_{k=1}^{\min(j,n^{\prime})}d_{k}(-)^{k+1}D_{j-k}\in\mathbb{Z}_{+}.$ (46)
However, Eq. (45) cannot be defined in $x\in(0,+\infty)$ as this would result to divergent and negative partition functions in $z_{-R}(x)$. Rather, this definition applies only to $x\in(0,x_{0})$, with $x_{0}=e^{-\beta\epsilon_{0}}$ under the following considerations:
  • Because $d_{0}=1$, i.e., there is only one vacuum state, we require that $1>|\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}|$ for the geometric series expansion to hold. However, because
    $z_{-R}(x)=\frac{1}{1+\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}}>0,$ (47)
    then it is automatically guaranteed that $1>-\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}$ is always true for appropriate values of $x$.
  • Consider otherwise that $x_{0}\to\infty$. In this case, we have
    $\lim_{x\to\infty}\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}=\begin{cases}+\infty,\quad\text{for even }n^{\prime}\\ -\infty,\quad\text{for odd }n^{\prime}\end{cases}.$ (48)
    However, as implied in the previous bullet, we require $\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}<1$ to be true within $x\in(0,x_{0}).$ Clearly, if $x_{0}\to\infty$, this condition is violated. Hence, we must impose $x_{0}$ to be finite, and therefore so must $\beta\epsilon_{0}$.
  • At the boundaries of $x\in(0,x_{0}),$ $z_{-R}(x)$ ceases to be of finite and positive value. However, $z_{-R}(x_{0})=1/z_{R}(-x_{0})\neq 0$ since $z_{R}(x)$ is a finite degree polynomial with positive finite coefficients, and therefore does not diverge for finite $x$. This therefore implies that
    $\lim_{x\to x_{0}^{-}}z_{-R}(x)=\lim_{\beta\epsilon\to\beta\epsilon_{0}^{+}}z_{-R}(e^{-\beta\epsilon})=+\infty,$ (49)
    meaning $x_{0}$ must be the smallest positive real root of $f(x)=z_{R}(-x)$.
Going back to the average occupation number definition in Eq. (7), we then have
$\langle\tilde{n}\rangle_{-R}=\frac{\sum_{n=1}^{n^{\prime}}nd_{n}(-)^{n-1}x^{n}}{\sum_{n=0}^{n^{\prime}}d_{n}(-x)^{n}}=\frac{x\partial_{-x}z_{R}(-x)}{z_{R}(-x)}.$ (50)
From the earlier discussions, we can therefore rewrite $z_{R}(-x)=g(-x)(-x+x_{0})^{\tau},$ with $\tau$ the multiplicity of the root $x_{0},$ and g(x)g(-x) a finite degree polynomial possibly containing other real roots larger than $x_{0}$. Then
$\langle\tilde{n}\rangle_{-R}$ $=x\frac{\partial_{-x}z_{R}(-x)}{z_{R}(-x)}$
$=x\frac{g^{\prime}(-x)(-x+x_{0})^{\tau}+\tau g(-x)(-x+x_{0})^{\tau-1}}{g(-x)(-x+x_{0})^{\tau}}$
$=\frac{xg^{\prime}(-x)}{g(-x)}+\frac{\tau x}{(-x+x_{0})},$ (51)
and hence $\langle\tilde{n}\rangle_{-R}$ also diverges at $x_{0}$. At the same time, because g(x)g(-x) has no real roots in $(0,x_{0}),$ then $\langle\tilde{n}\rangle_{-R}$ is smooth in this region. Finally, one can show that
$\lim_{\beta\epsilon\to\infty}\langle\tilde{n}\rangle_{-R}$ $=0,$ (52a)
@@
$\partial_{\beta\epsilon}\langle\tilde{n}\rangle_{-R}\Bigl|_{\beta\epsilon\to\infty}$ $\propto-d_{1},$ (52b)
@@
hence within the domain $\epsilon\in(\epsilon_{0},\infty)$, the average occupation number is exponentially decreasing from infinity, reminiscent of bose statistics as in Fig. 1(a). This motivates us to classify paraparticles that have single-mode partition functions in the form of Eq. (45) with $D_{n}\neq 0$ for all $n$ as $R$-parabosons.
@@
While the earlier discussion shows that all $R$-parafermions with a four-tensor $R$ have a corresponding $R$-parabosons described by R-R, the converse is not always true. Consider the examples
$M_{5}$ $=\begin{pmatrix}\zeta&0&0&0\\ 0&0&\kappa&0\\ 0&\kappa&0&0\\ \eta&0&0&-\zeta\end{pmatrix}$, (53a)
$M_{6}$ $=\begin{pmatrix}\zeta&0&0&\eta\\ 0&0&\kappa&0\\ 0&\kappa&0&0\\ 0&0&0&-\zeta\end{pmatrix}$, (53b)
@@
$M_{7}$ $=\begin{pmatrix}\zeta&0&0&0\\ 0&0&\delta&0\\ 0&1/\delta&0&0\\ 0&0&0&-\zeta\end{pmatrix}$, (53c)
where $\zeta,\kappa=\pm 1,\,\eta\in\mathbb{C},\,\delta\in\mathbb{C}_{\neq 0}$. In all cases, no exclusion principle can be extracted from their GCRs, and therefore all possible combinations of $\delta,\zeta,\eta,$ and $\kappa$ would result to $d_{n}\neq 0$ for all $n$. Their single-mode partition functions and average occupation number is therefore fully described by
$z(x)$ $=\sum_{n=0}^{\infty}d_{n}x^{n},$ (54a)
$\langle\tilde{n}\rangle$ $=\frac{xz^{\prime}(x)}{z(x)}=\frac{\sum_{n=1}^{\infty}nd_{n}x^{n}}{\sum_{n=0}^{\infty}d_{n}x^{n}}.$ (54b)
However, because $d_{n}\in\mathbb{Z}_{+}$, $z(x)$ can only admit positive real values up to some finite $x^{\prime}\in(0,1]$ such that
$\lim_{x\to(x^{\prime})^{-}}z(x)=\lim_{x\to x^{\prime}}\biggl(\sum_{n=0}^{\infty}d_{n}x^{n}\biggr)=+\infty.$ (55a)
This also results to
$\lim_{x\to(x^{\prime})^{-}}z^{\prime}(x)=\lim_{x\to x^{\prime}}\biggl(\sum_{n=1}^{\infty}nd_{n}x^{n-1}\biggr)=+\infty,$ (55b)
but because the power series structure of $z^{\prime}(x)$ grows faster than $z(x)$ for increasing index $n$, then
$\lim_{x\to(x^{\prime})^{-}}\frac{z^{\prime}(x)}{z(x)}=+\infty,$ (56)
and therefore $\langle\tilde{n}\rangle$ also diverges at $x^{\prime}$. Combined with the fact that
$\lim_{\beta\epsilon\to+\infty}\langle\tilde{n}\rangle=\lim_{x\to 0}\frac{xz^{\prime}(x)}{z(x)}=0,$ (57)
and with $z(x)$ and $z^{\prime}(x)$ being continuous and smooth between $x\in(0,x^{\prime})$, then for $x^{\prime}=e^{-\beta\epsilon_{0}}$, $\langle n\rangle_{R}$ should still be under the $R$-paraboson classification as in Fig. 1(a). Remarkably, this implies that for $R$-paraparticles, there are more $R$-parabosons than $R$-parafermions modulo internal basis transformations.